# On Mirzakhani’s twist torus conjecture

### v1

@@ 1. Introduction
$dp*\mu_{\operatorname{Mirz}}(X)={B(X)}~{}d\operatorname{vol}_{\operatorname{WP}}(X),$
@@
    if no component of $S\setminus\gamma$ has ()($\nabla$)( ∇ ), then $\mu_{\infty}$ is $\mu_{\operatorname{Mirz}}/b_{g}$.
@@
The limiting measure $\mu_{\infty}$ always has more structure than just earthquake-flow invariance. In [CF21], the authors extended the earthquake flow to a (measurable) action by “stretchquakes.” There and in [CF24], the authors showed that stretchquake-invariant probabilty measures on $\mathcal{P}^{1}\mathcal{M}_{g}$ correspond to Lebesgue-class measures on affine $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant subvarieties of strata of holomorphic quadratic differentials (§4.2). The measure $\mu_{\infty}$ is always stretchquake-invariant, and most of the time, we can identify it as corresponding to the Masur–Smillie–Veech measure on a component of a stratum of quadratic differentials with some zeros of order 2; see Corollary 5.12. We emphasize that Mirzakhani’s conjecture only references $\mu_{\operatorname{Mirz}}$; it seems to have been unexpected that families of twist tori could equidistribute to anything else.
@@
The strategy of our proof is to exploit a deep connection between the earthquake and horocycle flows discovered by Mirzakhani [Mir08] and further developed by the authors in [CF21, CF24]. Briefly, there is a Borel isomorphism $\mathcal{O}:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ that takes the earthquake flow to the horocycle flow on the moduli space $\mathcal{Q}^{1}\mathcal{M}_{g}$ of unit area quadratic differentials in a time-preserving way. While $\mathcal{O}$ is not continuous, it is continuous along the leaves of a certain foliation and each $\mathbb{T}_{\gamma}(\bm{\ell},\mathbf{h})$ is contained in such a leaf. This allows us to map the twist torus via $\mathcal{O}$ to get a torus in $\mathcal{Q}^{1}\mathcal{M}_{g}$ consisting of flat surfaces glued together from horizontal cylinders corresponding to the components of $\gamma$ with heights $\mathbf{h}$ and lengths $\bm{\ell}$; see Figure 1. Once in this setting, we can use powerful results from Teichmüller dynamics to understand the limiting distribution of expanding twist tori in $\mathcal{Q}^{1}\mathcal{M}_{g}$.
The driving engine of our proof is then the following theorem, which allows us to pull back equidistribution statements from $\mathcal{Q}^{1}\mathcal{M}_{g}$ to $\mathcal{P}^{1}\mathcal{M}_{g}$. This is a special case of [CF24, Theorem B].
@@ 1.1. Context: unipotent flows on moduli spaces
There are surprising similarities between the dynamics of groups generated by unipotents on homogeneous spaces and the action of $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$ on strata of holomorphic differentials. By way of [Mir08] and [CF21, CF24], parts of this analogy can now be extended to the action (of the group $P$ of upper triangular matrices in $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$) by stretchquakes on $\mathcal{P}^{1}\mathcal{M}_{g}$.
@@
Leveraging the ergodicity of the earthquake flow, Mirzakhani showed that the analogues of expanding horospheres ($L$-level sets of the hyperbolic length of a simple closed curve) equidistribute to $\pi_{*}\mu_{\operatorname{Mirz}}$ on $\mathcal{M}_{g}$ as $L\rightarrow\infty$ [Mir07]. These level sets can be lifted to $\mathcal{P}^{1}\mathcal{M}_{g}$ in a natural way by keeping track of the curve, as in §3.1. There are finite, earthquake flow–invariant measures on these lifted level sets that are uniformly expanded/contracted in forward/backward time by the “dilation flow,” the diagonal part of the stretchquake action (see the end of §2.3). The lifted level sets have the maximum dimension with this property; see also Theorems 2.1 and 2.2 that explain that the conjugacy $\mathcal{O}$ maps lifted level sets to leaves of the strong unstable foliation for the Teichmüller geodesic flow. These observations justify Mirzakhani’s use of the “horospherical” terminology in [Mir07] (compare [Wri20, §13.3]).
@@
As the dilation flow $\left(\operatorname{Dil}_{t}\right)$ normalizes the earthquake flow, pushforwards of horospherical (slice) measures by the dilation flow are themselves horospherical (slice) measures. Applying a theorem of Forni [For21] (see also [EMM22]), we arrive at the following:
@@
In the second case, the limiting measure is the pushforward via $\mathcal{O}^{-1}$ of the Masur–Smillie–Veech measure on the stratum component $\mathcal{Q}$ of $\mathcal{Q}^{1}\mathcal{M}_{g}$ corresponding to the horospherical slice; see §4.3 for the proof. The generic pair $(X,\lambda)$ with respect to one of these measures is a hyperbolic surface equipped with a measured lamination that cuts it into a union of regular ideal polygons [CF21, Corollary 1.2].
@@ 2.1. The horocycle flow
Strata are complex orbifolds of dimension $(2g-2)+n+\delta$, where $\delta=1$ if $\varepsilon=+1$ and $\delta=0$ otherwise. This can be seen via local period coordinates, constructed as follows. Every $q$ has a a holonomy double cover $\widehat{q}\to q$, branched over the zeros of odd order, so that $\widehat{q}$ is a (possibly disconnected) abelian differential with a holomorphic covering involution. Let $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ denote the $(-1)$-subspace for this involution, where $(\widehat{S},\widehat{Z})\cong(\widehat{q},Z(\widehat{q}))$. Take a triangulation of $q$ by saddle connections, singularity-free geodesics joining the cone points of $q$; this can be lifted to a triangulation $\widehat{\mathsf{T}}$ of $\widehat{q}$ by saddle connections. Integrating $\widehat{q}$ over the edges of $\widehat{\mathsf{T}}$ defines a cohomology class in $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$, and any small deformation of all of the triangles subject to the constraints imposed by $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ defines a new quadratic differential in the same stratum. The period vector mapping into $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ near $q$ is called a period coordinate chart. Two overlapping period coordinate charts differ by an integer linear transformation.
@@
There is an action of $\mathsf{SL}_{2}\mathbb{R}$ on $\mathcal{Q}^{1}\mathcal{M}_{g}$ preserving each stratum component $\mathcal{Q}$. For $M\in\mathsf{SL}_{2}\mathbb{R}$ and $q\in\mathcal{Q}$, $M\cdot q\in\mathcal{Q}$ is defined by postcomposing charts for the singular flat metric obtained by integration with $M:\mathbb{C}\to\mathbb{C}$. Since these charts are only defined up to ±$\pm 1$, this $\mathsf{SL}_{2}\mathbb{R}$ action factors through $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$. In a period coordinate chart $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}\cong\mathbb{C}\otimes H^{1}(\widehat{S},\widehat{Z};\mathbb{R})^{-}$, the action is given by applying $M$ to $\mathbb{C}$. The one-parameter subgroups
@@
There are strong stable and strong unstable foliations for the Teichmüller geodesic flow. At every point $q$ the strong stable (unstable) leaf through $q$ locally consists of those $q^{\prime}$ with the same real (imaginary) foliation as $q$. Each unstable leaf
@@
therefore has a stratified real-analytic structure coming from its intersections with strata of $\mathcal{QT}_{g}$. For every $\lambda$, define $\mathcal{MF}(\lambda):=\{\eta\in\mathcal{MF}_{g}\mid(\eta,\lambda)\notin\Delta\}.$ Using (1), we have
@@
which also endows $\mathcal{MF}(\lambda)$ with a stratified real-analytic structure.
@@ 2.2. The earthquake flow
coincides with the imaginary part of the Weil–Petersson Kähler form, hence is invariant by $\operatorname{Mod}_{g}$ [Wol85]. The top wedge power of $\omega_{\operatorname{WP}}$ (suitably normalized) defines a volume form and corresponding finite Weil–Petersson measure $\operatorname{vol}_{\operatorname{WP}}$ in the class of Lebesgue on $\mathcal{M}_{g}$.
@@ 2.3. The conjugacy
There is an important connection between the horocycle and earthquake flows introduced above: they are Borel isomorphic (see Theorem 2.2 below). This relationship was first observed by Mirzakhani on the full $\mu_{\operatorname{Mirz}}$-measure subset of pairs $(X,\lambda)\in\mathcal{P}^{1}\mathcal{M}_{g}$ where $\lambda$ is a maximal geodesic lamination [Mir08] and was extended to the entirety of $\mathcal{P}^{1}\mathcal{M}_{g}$ by the authors [CF21].
@@
Because their leaves are always orthogonal to $\lambda$, the orthogeodesic foliations on the components of $X\setminus\lambda$ can be extended continuously across $\lambda$, yielding a singular foliation $\mathcal{O}_{\lambda}(X)$ on $S$. This foliation also carries a natural transverse measure, defined by assigning to any geodesic arc in $\lambda$ its Lebesgue measure and then extending uniquely to all arcs via transverse isotopy. See [CF21, Section 5] for more details.
This defines a map $\mathcal{O}_{\lambda}:\mathcal{T}_{g}\to\mathcal{MF}_{g}$ recording the equivalence class of $\mathcal{O}_{\lambda}(X)$; the following theorem records the global properties of $\operatorname{\mathcal{O}_{\lambda}}$. When $\lambda$ is maximal, it is due to Thurston [Thu98]; see also [Bon96] for a treatment in terms of transverse Hölder distributions. For arbitrary $\lambda$, this is Theorem E of [CF21].
@@
Moreover, for $(X,\lambda)\in\mathcal{PT}_{g}$, the real foliation of $\mathcal{O}(X,\lambda)$ is actually isotopic to $\mathcal{O}_{\lambda}(X)$ (see [CF21, §5.3]).
@@
of the same name. The following omnibus theorem combines the main theorems of [Mir08] and [CF21] together with a special case of the main result of [CF24].
@@
The map $\mathcal{O}:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ is a bijective Borel isomorphism with the following properties:
@@
    $\mathcal{O}$ takes earthquake flow to horocycle flow in a time-preserving way: $\mathcal{O}\circ\operatorname{Eq}_{s}=u_{s}\circ\mathcal{O}$.
@@
    $\mathcal{O}_{*}\mu_{\operatorname{Mirz}}/b_{g}$ is the Masur–Smillie–Veech probability measure on the principal stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$.
@@
    Each $k$-pronged singularity of $\operatorname{\mathcal{O}_{\lambda}}(X)$ corresponds to an order $(k-2)$ zero of $\mathcal{O}(X,\lambda)$ and vice versa.
@@
    If $\mathcal{Q}$ is a component of a stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$ and $q\in\mathcal{Q}$ has no horizontal saddle connections, then $q$ is a point of continuity for $\mathcal{O}^{-1}|_{\mathcal{Q}}$.
@@
Mirzakhani pointed out that $\mathcal{O}$ (or any extension of her original conjugacy) is discontinuous along sequences $(X_{n},\lambda_{n})$ where the measured and Hausdorff limits of maximal laminations $\lambda_{n}$ disagree [Mir08, p. 33]. It is now known that there can be no continuous map $\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ conjugating the earthquake and horocycle flows [AHW23], so Theorem 2.2 (and the more general Theorem A of [CF24]) are essentially optimal.
The conjugacy $\mathcal{O}$ allows us to pull back the Teichmüller geodesic flow from $\mathcal{Q}^{1}\mathcal{M}_{g}$ to define a measurable dilation flow $(\operatorname{Dil}_{t})$ on $\mathcal{P}^{1}\mathcal{M}_{g}$ that normalizes the earthquake flow (see [CF21, §15.3]):

### v2

@@ 1. Introduction
$dp_{*}\mu_{\operatorname{Mirz}}(X)={B(X)}~{}d\operatorname{vol}_{\operatorname{WP}}(X),$
@@
    if $g\geq 3$ and no component of $S\setminus\gamma$ has ()($\nabla$)( ∇ ), then $\mu_{\infty}$ is $\mu_{\operatorname{Mirz}}/b_{g}$.
@@
Remark 1.3.
For $g=2$, statement (1) of Theorem 1.2 would follow, using our techniques, from [AW23, Conjecture 1.2] about “almost high rank” invariant submanifolds; see §§45.
@@
The limiting measure $\mu_{\infty}$ always has more structure than just earthquake-flow invariance. In [CF24b], the authors extended the earthquake flow to a (measurable) action by “stretchquakes.” There and in [CF24a], the authors showed that stretchquake-invariant probabilty measures on $\mathcal{P}^{1}\mathcal{M}_{g}$ correspond to Lebesgue-class measures on affine $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant subvarieties of strata of holomorphic quadratic differentials (§4.2). The measure $\mu_{\infty}$ is always stretchquake-invariant, and most of the time, we can identify it as corresponding to the Masur–Smillie–Veech measure on a component of a stratum of quadratic differentials with some zeros of order 2; see Corollary 5.12. We emphasize that Mirzakhani’s conjecture only references $\mu_{\operatorname{Mirz}}$; it seems to have been unexpected that families of twist tori could equidistribute to anything else.
@@
The strategy of our proof is to exploit a deep connection between the earthquake and horocycle flows discovered by Mirzakhani [Mir08] and further developed by the authors in [CF24b, CF24a]. Briefly, there is a Borel isomorphism $\mathcal{O}:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ that takes the earthquake flow to the horocycle flow on the moduli space $\mathcal{Q}^{1}\mathcal{M}_{g}$ of unit area quadratic differentials in a time-preserving way. While $\mathcal{O}$ is not continuous, it is continuous along the leaves of a certain foliation and each $\mathbb{T}_{\gamma}(\bm{\ell},\mathbf{h})$ is contained in such a leaf. This allows us to map the twist torus via $\mathcal{O}$ to get a torus in $\mathcal{Q}^{1}\mathcal{M}_{g}$ consisting of flat surfaces glued together from horizontal cylinders corresponding to the components of $\gamma$ with heights $\mathbf{h}$ and lengths $\bm{\ell}$; see Figure 1. Once in this setting, we can use powerful results from Teichmüller dynamics to understand the limiting distribution of expanding twist tori in $\mathcal{Q}^{1}\mathcal{M}_{g}$.
The driving engine of our proof is then the following theorem, which allows us to pull back equidistribution statements from $\mathcal{Q}^{1}\mathcal{M}_{g}$ to $\mathcal{P}^{1}\mathcal{M}_{g}$. This is a special case of [CF24a, Theorem B].
@@ 1.1. Context: unipotent flows on moduli spaces
There are surprising similarities between the dynamics of groups generated by unipotents on homogeneous spaces and the action of $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$ on strata of holomorphic differentials. By way of [Mir08] and [CF24b, CF24a], parts of this analogy can now be extended to the action (of the group $P$ of upper triangular matrices in $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$) by stretchquakes on $\mathcal{P}^{1}\mathcal{M}_{g}$.
@@
Leveraging the ergodicity of the earthquake flow, Mirzakhani showed that the analogues of expanding horospheres ($L$-level sets of the hyperbolic length of a simple closed curve) equidistribute to $\pi_{*}\mu_{\operatorname{Mirz}}$ on $\mathcal{M}_{g}$ as $L\rightarrow\infty$ [Mir07]. These level sets can be lifted to $\mathcal{P}^{1}\mathcal{M}_{g}$ in a natural way by keeping track of the curve, as in §3.1. There are finite, earthquake flow–invariant measures on these lifted level sets that are uniformly expanded/contracted in forward/backward time by the “dilation flow,” the diagonal part of the stretchquake action (see the end of §2.3). The lifted level sets have the maximum dimension with this property; see also Theorems 2.1 and 2.3 that explain that the conjugacy $\mathcal{O}$ maps lifted level sets to leaves of the strong unstable foliation for the Teichmüller geodesic flow. These observations justify Mirzakhani’s use of the “horospherical” terminology in [Mir07] (compare [Wri20, §13.3]).
@@
As the dilation flow $\left(\operatorname{Dil}_{t}\right)$ normalizes the earthquake flow, pushforwards of horospherical (slice) measures by the dilation flow are themselves horospherical (slice) measures. Applying a result of Forni [For21, Theorem 1.6] (see also [EMM22, Proposition 3.2]), we arrive at the following:
@@
In the second case, the limiting measure is the pushforward via $\mathcal{O}^{-1}$ of the Masur–Smillie–Veech measure on the stratum component $\mathcal{Q}$ of $\mathcal{Q}^{1}\mathcal{M}_{g}$ corresponding to the horospherical slice; see §4.3 for the proof. The generic pair $(X,\lambda)$ with respect to one of these measures is a hyperbolic surface equipped with a measured lamination that cuts it into a union of regular ideal polygons [CF24b, Corollary 1.2].
@@ 2.1. The horocycle flow
Strata are complex orbifolds of dimension $(2g-2)+n+\delta$, where $\delta=1$ if $\varepsilon=+1$ and $\delta=0$ otherwise. This can be seen via local period coordinates, constructed as follows. Every $q$ has a a holonomy double cover $\widehat{q}\to q$, branched over the zeros of odd order, so that $\widehat{q}$ is a (possibly disconnected) abelian differential with a holomorphic covering involution. Let $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ denote the $(-1)$-subspace for this involution, where $(\widehat{S},\widehat{Z})\cong(\widehat{q},Z(\widehat{q}))$. Take a triangulation of $q$ by saddle connections, singularity-free geodesics joining the cone points of $q$; this can be lifted to a triangulation $\widehat{\mathsf{T}}$ of $\widehat{q}$ by saddle connections. Integrating $\widehat{q}$ over the edges of $\widehat{\mathsf{T}}$ defines a cohomology class in $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$, and any small deformation of all of the triangles subject to the constraints imposed by $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ defines a new quadratic differential in the same stratum. The period vector mapping into $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ near $q$ is called a period coordinate chart. Two overlapping period coordinate charts differ by an integer linear transformation. We refer the reader to [Zor06], [Wri20], and [AM24] for a more thorough discussion of period coordinates and background on flat surfaces.
@@
There is an action of $\mathsf{SL}_{2}\mathbb{R}$ on $\mathcal{Q}^{1}\mathcal{M}_{g}$ preserving each stratum component $\mathcal{Q}$. For $M\in\mathsf{SL}_{2}\mathbb{R}$ and $q\in\mathcal{Q}$, $M\cdot q\in\mathcal{Q}$ is defined by postcomposing charts for the singular flat metric obtained by integration with the $\mathbb{R}$-linear map $M:\mathbb{C}\to\mathbb{C}$. Since these charts are only defined up to ±$\pm 1$, this $\mathsf{SL}_{2}\mathbb{R}$ action factors through $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$. In a period coordinate chart $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}\cong\mathbb{C}\otimes H^{1}(\widehat{S},\widehat{Z};\mathbb{R})^{-}$, the action is given by applying $M$ to $\mathbb{C}$. The one-parameter subgroups
@@
Following [Vee86] and [For02] (see also [ABEM12, Theorem 3.15]), we define the strong stable and strong unstable foliations for the Teichmüller geodesic flow as follows. At every point $q$ the strong stable (unstable) leaf through $q$ locally consists of those $q^{\prime}$ with the same real (imaginary) foliation as $q$. Each unstable leaf
@@
therefore has a stratified real-analytic structure coming from restricting the linear structure of strata of $\mathcal{QT}_{g}$ to intersections with $\mathcal{W}^{uu}(\lambda)$. For every $\lambda$, define $\mathcal{MF}(\lambda):=\{\eta\in\mathcal{MF}_{g}\mid(\eta,\lambda)\notin\Delta\}.$ Using (1), we have
@@
which we can use to equip $\mathcal{MF}(\lambda)$ with a stratified real-analytic structure.
@@ 2.2. The earthquake flow
coincides with the imaginary part of the Weil–Petersson Kähler form, hence is invariant by $\operatorname{Mod}_{g}$ [Wol85]. The top wedge power of $\omega_{\operatorname{WP}}$ defines a volume form and corresponding Weil–Petersson measure $\operatorname{vol}_{\operatorname{WP}}$ on $\mathcal{M}_{g}$. This measure is in the class of Lebesgue and has finite total mass.
@@ 2.3. The conjugacy
There is an important connection between the horocycle and earthquake flows introduced above: they are Borel isomorphic (see Theorem 2.3 below). This relationship was first observed by Mirzakhani on the full $\mu_{\operatorname{Mirz}}$-measure subset of pairs $(X,\lambda)\in\mathcal{P}^{1}\mathcal{M}_{g}$ where $\lambda$ is a maximal geodesic lamination [Mir08] and was extended to the entirety of $\mathcal{P}^{1}\mathcal{M}_{g}$ by the authors [CF24b].
@@
Because their leaves are always orthogonal to $\lambda$, the orthogeodesic foliations on the components of $X\setminus\lambda$ can be extended continuously across $\lambda$, yielding a singular foliation $\mathcal{O}_{\lambda}(X)$ on $S$. This foliation also carries a natural transverse measure, defined by assigning to any geodesic arc in $\lambda$ its Lebesgue measure and then extending uniquely to all arcs via transverse isotopy. See [CF24b, Section 5] for more details.
This defines a map $\mathcal{O}_{\lambda}:\mathcal{T}_{g}\to\mathcal{MF}_{g}$ recording the equivalence class of $\mathcal{O}_{\lambda}(X)$; the following theorem records the global properties of $\operatorname{\mathcal{O}_{\lambda}}$. When $\lambda$ is maximal, it is due to Thurston [Thu98]; see also [Bon96] for a treatment in terms of transverse Hölder distributions. For arbitrary $\lambda$, the following is a direct consequence of Theorems D and E of [CF24b].
@@
Remark 2.2.
Theorem D of [CF24b] establishes that $\operatorname{\mathcal{O}_{\lambda}}$ is a homeomorphism. Theorem E therein asserts that $\operatorname{\mathcal{O}_{\lambda}}$ factors as $\operatorname{\mathcal{O}_{\lambda}}=I_{\lambda}^{-1}\circ\sigma_{\lambda}$, where $\sigma_{\lambda}$ is stratified real-analytic and $I_{\lambda}^{-1}$ is linear when restricted to each stratum. Thus $\operatorname{\mathcal{O}_{\lambda}}$ is also stratified real-analytic.
@@
Moreover, for $(X,\lambda)\in\mathcal{PT}_{g}$, the real foliation of $\mathcal{O}(X,\lambda)$ is actually isotopic to $\mathcal{O}_{\lambda}(X)$ (see [CF24b, §5.3]).
@@
of the same name. The following omnibus theorem combines the main theorems of [Mir08] and [CF24b] together with a special case of the main result of [CF24a].
@@
The map $\mathcal{O}:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ is a bijective Borel isomorphism [^3] Theorem C of [CF24b] and Theorem 15.4 of [CF24a]. with the following properties:
@@
    $\mathcal{O}$ takes earthquake flow to horocycle flow in a time-preserving way: $\mathcal{O}\circ\operatorname{Eq}_{s}=u_{s}\circ\mathcal{O}$ [^4] Theorem A of [CF24b]; see also Theorem 1.1 of [Mir08].
@@
    $\mathcal{O}_{*}\mu_{\operatorname{Mirz}}/b_{g}$ is the Masur–Smillie–Veech probability measure on the principal stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$. [^5] Section 1.2 of [Mir08].
@@
    Each $k$-pronged singularity of $\operatorname{\mathcal{O}_{\lambda}}(X)$ corresponds to an order $(k-2)$ zero of $\mathcal{O}(X,\lambda)$ and vice versa. [^6] Proposition 5.10 of [CF24b].
@@
    If $\mathcal{Q}$ is a component of a stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$ and $q\in\mathcal{Q}$ has no horizontal saddle connections, then $q$ is a point of continuity for $\mathcal{O}^{-1}|_{\mathcal{Q}}$. [^7] A stronger version of this is proved as Theorem 15.6 of [CF24a].
@@
Mirzakhani pointed out that $\mathcal{O}$ (or any extension of her original conjugacy) is discontinuous along sequences $(X_{n},\lambda_{n})$ where the measured and Hausdorff limits of maximal laminations $\lambda_{n}$ disagree [Mir08, p. 33]. It is now known that there can be no continuous map $\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ conjugating the earthquake and horocycle flows [AHW23], so Theorem 2.3 (and the more general Theorem A of [CF24a]) are essentially optimal.
The conjugacy $\mathcal{O}$ allows us to pull back the Teichmüller geodesic flow from $\mathcal{Q}^{1}\mathcal{M}_{g}$ to define a measurable dilation flow $(\operatorname{Dil}_{t})$ on $\mathcal{P}^{1}\mathcal{M}_{g}$ that normalizes the earthquake flow (see [CF24b, §15.3]):